# Agent Lumos: Unified and Modular Training for Open-Source Language Agents

### v2

@@ Abstract
To foster generalizable agent learning, we collect large-scale, unified, and high-quality training annotations derived from diverse ground-truth reasoning rationales across various complex interactive tasks. On 9 datasets, Lumos exhibits several key advantages: (1) Lumos excels multiple larger open-source agents on the held-out datasets (unused for training) for each task type. Lumos even surpasses GPT agents on QA and web tasks; (2) Lumos outperforms open-source agents produced by chain-of-thoughts and unmodularized integrated training; and (3) Lumos effectively generalizes to unseen tasks, outperforming 33B-scale agents and domain-specific agents. Code and data will be released.
@@
Da Yin††thanks: Work was done during Da’s internship at AI2. Code: https://github.com/allenai/lumos. Models & Data: https://huggingface.co/ai2lumos Faeze Brahman Abhilasha Ravichander Khyathi Chandu Kai-Wei Chang Yejin Choi Bill Yuchen Lin Allen Institute for AI UCLA University of Washington https://allenai.github.io/lumos/ da.yin@ucla.edu yuchenl@allenai.org
@@ 1 Introduction
Competitive Performance.
Lumos outperforms a great number of open-source agents on the Lumos held-out datasets unused in Lumos training data across the four training task types. Lumos even surpasses GPT-based agents in web and QA tasks. Specifically, Lumos shows a 5.0% enhancement over GPT-4 on Mind2Web, and 4.1% and 3.5% LLM accuracy111A metric defined in Xu et al. (2023) to identify the semantic equivalence between predictions and gold answers. improvement on HotpotQA over the ReAct and ReWOO agents fully based on GPT-3.5-turbo, respectively.
@@
We evaluate Lumos on two unseen tasks, WebShop (Yao et al., 2022a), a text game for online shopping, and InterCode Yang et al. (2023), an interactive code generation task. Lumos even surpasses 30B-scale agents, especially by nearly 20 reward points on WebShop. Lumos also delivers a consistent reward improvement over domain-specific agents. This suggests that Lumos can generalize across tasks, hinting at potential benefits for a wide spectrum of language agent applications.
@@ 4.1 Experimental Setup
| AgentLM-70B | 13.5 |
| Lumos-I | 27.6 |
| Lumos-I-13B | 31.3 |
@@
| Lumos-O | 50.5 | 65.5 |
| Lumos-I | 47.1 | 63.6 |
| Lumos-O-13B | 55.4 | 69.3 |
@@
| Lumos-O | 70.1 | 56.9 |
| Lumos-I | 71.3 | 58.4 |
| Lumos-I-13B | 72.4 | 58.2 |
@@
| ReWOO-open | LLAMA-7B | GPT-3.5-turbo | 56 | 37 / - |
@@
| FiReAct | LLAMA-2-7B | LLAMA-2-7B | - | - / 26.2 |
| FiReAct | CodeLLAMA-34B | CodeLLAMA-34B | - | - / 27.8 |
| Lumos-O | LLAMA-2-7B | GPT-3.5-turbo | 60.6 | 39.2 / 24.9 |
| Lumos-I | LLAMA-2-7B | LLAMA-2-7B | 58.3 | 37.3 / 23.5 |
| Lumos-I | LLAMA-2-7B | GPT-3.5-turbo | 65.7 | 45.9 / 29.4 |
| Lumos-I | LLAMA-2-7B | GPT-4 | 72.4 | 56.8 / 36.0 |
| Lumos-I | LLAMA-2-13B | GPT-3.5-turbo | 65.3 | 50.2 / 31.4 |
| Lumos-I | LLAMA-2-13B | GPT-4 | 76.7 | 57.4 / 36.3 |
@@
| Lumos-I-13B | 46.2 | 4.2 |
| Lumos-I-13B | 45.7 | 5.8 |
| Lumos-I-13B | 47.3 | 3.5 |
| Lumos-I-13B | 43.8 | 4.0 |
| Lumos-I-13B | 50.3 | 7.3 |
@@ 4.2 Training Task Performance
We evaluate Lumos across an array of complex interactive tasks - QA, web, math and multimodal tasks. The evaluation mainly follows the settings established by AgentBench (Liu et al., 2023b) and ReWOO (Xu et al., 2023) (see App. H). For each task type, excluding web tasks, we include a held-out dataset to assess the model’s generalizability across the domains within the same task category. The performance is displayed in Tab. 1. Note that in Tab. 1, task-specific agents Lumos-I are trained using task-specific data belonging to task type (e.g., , , , ). Lumos-I represents the agent after unified training with the combination of four task type annotations. More evaluation details are shown in App. H.
@@ 4.4 Lumos Generalizability Evaluation
| Lumos-I w/ Low-Level Subgoals | 63.3 | 44.3 |
| Lumos-I Data | 65.7 | 45.9 |
To examine the generalization ability of Lumos, we evaluate it on the unseen tasks, WebShop (Yao et al., 2022a) and InterCode. WebShop resembles a text game444WebShop utilizes four actions in its training annotations: Search, FeatureRetrieve, Pick, and Click. The argument of Click is a shopping item, differing from that of Click in Mind2Web which includes an HTML element description., with its shopping environment and action space considerably differing from those covered in the training annotations of Lumos. InterCode (Yang et al., 2023) is an interactive code generation task that requires the agent to generate SQL code based on the external databases and involves unseen SQL commands. To make Lumos adapt to these unseen tasks, we supplement 2-/3-shot examples in the module inputs, enabling them to learn how to generate subgoals and ground to new sets of available actions (see App. J).
@@
| Lumos-O | 50.5 | 65.5 |
| Lumos-I | 47.1 | 63.6 |
@@
| Lumos-O | 70.1 | 56.9 |
| Lumos-I | 71.3 | 58.4 |
As outlined in Tab. 0(e), Lumos-I achieves a 3-7 higher average reward than domain-specific agents on WebShop. It also significantly outperforms larger agents such as WizardLM-30B and Vicuna-v1.3-33B (Chiang et al., 2023). We observe the similar trend on InterCode. It suggests that unified training enables agents to enhance the generalization on unseen tasks and novel actions. We provide more unified training results in App. F to show unified training also boosts the performance on most training task types.
@@ 4.5 Further Analysis on Training Annotations
We aim to address two questions pertinent to quality and format decisions. Q1: How good are our converted training annotations? Q2: Would adopting low-level subgoals be more effective than using high-level subgoals?
@@ Appendix B Statistics of Converted Training Annotations
To train agents like Lumos-I mentioned in Tab. 0(b), we need to leverage the annotations converted from 19778 data specific to math domain. For training a unified agent such as Lumos-I, we would use the annotations transformed from all the listed data as training set.
@@ Appendix G Execution Tools Associated with Action Interfaces
For the unseen task InterCode, the action interfaces include all the possible commands and functions provided in SQL programming language.
@@ I.4 In-Context Example For Obtaining Multimodal Task Annotations
Please convert natural language plans into a series of subgoals, their corresponding actions that lead to the successful implementation with respect to the given instructions. When generating the actions, please also attach the action’s results contained in the natural language plans. Please use ’R[number]’ to represent the execution results for each action. Please also use functions to represent the corresponding actions. For the actions, they must be one of the available actions, ’QA’, ’VQA’. Example 1: Task: If the cameraman were driving what do they have to do from this position? There’re some choices: A. turn left, B. drive straight, C. reverse course, D. turn right. Natural language plan: The would have to turn right because the lane has right turn arrows painted on it. The arrow on the street indicates that this lane can only go in one direction at the intersection. The sign on the road says to turn right. Overall, the final answer is ’turn right’. Subgoal-based plan: Subgoal 1: Describe the shape of the sign on the road lane the cameraman is in from the image. Action 1-1: R1 = VQA([IMG], Question: What’s the sign on the road lane?) = There’s a right turn arrow on the road. Subgoal 2: Answer which lane the cameraman is in. Action 2-1: R2 = QA([R1], Question: Which lane is the cameraman in?) = The cameraman is in right turn lane. Subgoal 3: Answer what the cameraman has to do from his position. Action 3-1: R3 = QA([R1, R2], Question: What does the cameraman have to do from his position?) = turn right. Subgoal 4: Select the choice from the answer choices. Action 4-1: R4 = VQA([IMG], Question: If the cameraman were driving what do they have to do from this position? We can also see that There’s a right turn arrow on the road, The cameraman is in right turn lane, turn right. There’re some choices: A. turn left, B. drive straight, C. reverse course, D. turn right. You must choose one from ’A’, ’B’, ’C’ or ’D’. Output only one letter!) = D
@@
Appendix K In-Context Example for Adapting Modules to Unseen Task — InterCode
Similar to §J, as InterCode is another unseen task, we showcase one of the three in-context InterCode-related examples used for the inference of planning and grounding modules, respectively.
K.1 In-Context InterCode-Related Example for Planning Module
@@
K.2 In-Context InterCode-Related Example for Grounding Module

### v3

# ACL 2024 Main Conference Agent Lumos: Unified and Modular Training for Open-Source Language Agents

@@ Abstract
To foster generalizable agent learning, we collect large-scale, unified, and high-quality training annotations derived from diverse ground-truth reasoning rationales across various complex interactive tasks. On 9 datasets, Lumos exhibits several key advantages: (1) Lumos excels multiple larger open-source agents on the held-out datasets (unused for training) for each task type. Lumos even surpasses GPT agents on QA and web tasks; (2) Lumos outperforms open-source agents produced by chain-of-thoughts and unmodularized integrated training; and (3) Lumos effectively generalizes to unseen tasks, outperforming 33B-scale agents and domain-specific agents.
[ backgroundcolor=purple!10, skipabove=1em, skipbelow=0em, leftline=true, topline=false, bottomline=false, rightline=false, linecolor=purple!88, linewidth=4pt ]githubquote
ACL 2024 Main Conference
@@
Da Yin♡††thanks: Work was done during Da’s internship at AI2. Code: https://github.com/allenai/lumos. Models & Data: https://huggingface.co/ai2lumos Faeze Brahman♠ Abhilasha Ravichander♠ Khyathi Chandu♠ Kai-Wei Chang♡ Yejin Choi♠♢ Bill Yuchen Lin♠ ♠Allen Institute for AI ♡UCLA ♢University of Washington https://allenai.github.io/lumos/ da.yin@cs.ucla.edu yuchenl@allenai.org
@@ 1 Introduction
Competitive Performance Across Tasks and Agent Training Formulations.
Lumos outperforms a great number of open-source agents on the Lumos held-out datasets unused in Lumos training data across the four training task types. Lumos even surpasses GPT-based agents in web and QA tasks. Specifically, Lumos shows a 5.0% enhancement over GPT-4 on Mind2Web, and 4.1% and 3.5% LLM accuracy111A metric defined in Xu et al. (2023) to identify the semantic equivalence between predictions and gold answers. improvement on HotpotQA over the ReAct and ReWOO agents fully based on GPT-3.5-turbo, respectively. Furthermore, we observe that Lumos training formulations outperform other potential agent training methods, such as chain-of-thoughts and integrated training, which instruct a single module to both plan and ground.
@@
We evaluate Lumos on two unseen tasks, WebShop (Yao et al., 2022a), a text game for online shopping, and InterCodeSQL Yang et al. (2023), an interactive code generation task. Lumos even surpasses 30B-scale agents, especially by nearly 20 reward points on WebShop. Lumos also delivers a consistent reward improvement over domain-specific agents. This suggests that Lumos can generalize across tasks, hinting at potential benefits for a wide spectrum of language agent applications.
@@ 4.1 Experimental Setup
| AgentLM-70B | 13.5★ |
| Lumos-IWeb | 27.6★ |
| Lumos-IWeb-13B | 31.3★ |
@@
| Lumos-OMath | 50.5★ | 65.5 |
| Lumos-IMath | 47.1★ | 63.6 |
| Lumos-OMath-13B | 55.4★ | 69.3 |
@@
| Lumos-OMM | 70.1★ | 56.9 |
| Lumos-IMM | 71.3★ | 58.4 |
| Lumos-IMM-13B | 72.4★ | 58.2 |
@@
| ReWOO-open | LLAMA-7B | GPT-3.5-turbo | 56 | 37 / -★ |
@@
| FiReAct | LLAMA-2-7B | LLAMA-2-7B | - | - / 26.2★ |
| FiReAct | CodeLLAMA-34B | CodeLLAMA-34B | - | - / 27.8★ |
| Lumos-OQA | LLAMA-2-7B | GPT-3.5-turbo | 60.6★ | 39.2 / 24.9 |
| Lumos-IQA | LLAMA-2-7B | LLAMA-2-7B | 58.3★ | 37.3 / 23.5 |
| Lumos-IQA | LLAMA-2-7B | GPT-3.5-turbo | 65.7★ | 45.9 / 29.4 |
| Lumos-IQA | LLAMA-2-7B | GPT-4 | 72.4★ | 56.8 / 36.0 |
| Lumos-IQA | LLAMA-2-13B | GPT-3.5-turbo | 65.3★ | 50.2 / 31.4 |
| Lumos-IQA | LLAMA-2-13B | GPT-4 | 76.7★ | 57.4 / 36.3 |
@@
| Lumos-IWeb-13B | 46.2 | 4.2 |
| Lumos-IMath-13B | 45.7 | 5.8 |
| Lumos-IQA-13B | 47.3 | 3.5 |
| Lumos-IMM-13B | 43.8 | 4.0 |
| Lumos-IAll-13B | 50.3 | 7.3 |
@@ 4.2 Training Task Performance
We evaluate Lumos across an array of complex interactive tasks - QA, web, math and multimodal tasks. The evaluation mainly follows the settings established by AgentBench (Liu et al., 2023b) and ReWOO (Xu et al., 2023) (see App. H). For each task type, excluding web tasks, we include a held-out dataset to assess the model’s generalizability across the domains within the same task category. The performance is displayed in Tab. 1. Note that in Tab. 1, task-specific agents Lumos-IX are trained using task-specific data belonging to task type (e.g., , , , ). Lumos-IAll represents the agent after unified training with the combination of four task type annotations. More evaluation details are shown in App. H.
@@ 4.4 Lumos Generalizability Evaluation
| Lumos-IQA w/ Low-Level Subgoals | 63.3 | 44.3 |
| Lumos-IQA Data | 65.7 | 45.9 |
To examine the generalization ability of Lumos, we evaluate it on the unseen tasks, WebShop (Yao et al., 2022a) and InterCodeSQL. WebShop resembles a text game444WebShop utilizes four actions in its training annotations: Search, FeatureRetrieve, Pick, and Click. The argument of Click is a shopping item, differing from that of Click in Mind2Web which includes an HTML element description., with its shopping environment and action space considerably differing from those covered in the training annotations of Lumos. InterCodeSQL (Yang et al., 2023) is an interactive code generation task that requires the agent to generate SQL code based on the external databases and involves unseen SQL commands. To make Lumos adapt to these unseen tasks, we supplement 2-/3-shot examples in the module inputs, enabling them to learn how to generate subgoals and ground to new sets of available actions (see App. J).
@@
| Lumos-OMath | 50.5 | 65.5 |
| Lumos-IMath | 47.1 | 63.6 |
@@
| Lumos-OMM | 70.1 | 56.9 |
| Lumos-IMM | 71.3 | 58.4 |
As outlined in Tab. 1(e), Lumos-I achieves a 3-7 higher average reward than domain-specific agents on WebShop. It also significantly outperforms larger agents such as WizardLM-30B and Vicuna-v1.3-33B (Chiang et al., 2023). We observe the similar trend on InterCodeSQL. It suggests that unified training enables agents to enhance the generalization on unseen tasks and novel actions. We provide more unified training results in App. F to show unified training also boosts the performance on most training task types.
@@ 4.5 Further Analysis on Training Annotations
We aim to address three questions pertinent to quality and format decisions. Q1: How good are our converted training annotations? Q2: Would adopting low-level subgoals be more effective than using high-level subgoals? Q3: Would incorporating Lumos annotation with general instruction-tuning data achieve comparable performance on instruction-following benchmarks?
@@
Effect on General Instruction-Following Benchmark.
As Lumos adopts the conversational style discussed in Sec. 3, any tasks that can be formalized as conversation tasks could be accommodated. Therefore, Lumos supports all the instruction-tuning tasks and complex interactive tasks. We conduct an experiment to show the effectiveness of Lumos in instruction tuning tasks. By integrating the general instruction-tuning training data, Alpaca, with the unified training annotations of Lumos, we created an expanded training dataset. To guide Lumos towards generating straightforward responses, we appended the phrase “Respond to me directly” to each input instance from Alpaca, instructing the model to avoid elaborating reasoning processes for the general instruction-tuning tasks. Tested on the Super-NaturalInstruction benchmark Wang et al. (2022), Lumos achieved a 39.3 ROUGE-L score, which is close to the LLAMA-2-7B trained on Alpaca alone (39.8), which suggesting the possibility of developing agent reasoning ability and general instruction following ability simultaneously.
@@ 6 Conclusion
Acknowledgement
This work was funded in part by DARPA MCS program through NIWC Pacific (N66001-19-2- 4031), DARPA ECOLE (#HR00112390060), and the Allen Institute for AI. We thank Mosaic team members, and the anonymous reviewers for the helpful discussions.
@@ Appendix B Statistics of Converted Training Annotations
To train agents like Lumos-IMath mentioned in Tab. 1(b), we need to leverage the annotations converted from 19778 data specific to math domain. For training a unified agent such as Lumos-I, we would use the annotations transformed from all the listed data as training set.
@@ Appendix G Execution Tools Associated with Action Interfaces
For the unseen task InterCodeSQL, the action interfaces include all the possible commands and functions provided in SQL programming language.
@@ I.4 In-Context Example For Obtaining Multimodal Task Annotations
Please convert natural language plans into a series of subgoals, their corresponding actions that lead to the successful implementation with respect to the given instructions. When generating the actions, please also attach the action’s results contained in the natural language plans. Please use ’R[number]’ to represent the execution results for each action. Please also use functions to represent the corresponding actions. For the actions, they must be one of the available actions, ’QA’, ’VQA’.
Example 1:
Task: If the cameraman were driving what do they have to do from this position? There’re some choices: A. turn left, B. drive straight, C. reverse course, D. turn right.
Natural language plan:
The would have to turn right because the lane has right turn arrows painted on it. The arrow on the street indicates that this lane can only go in one direction at the intersection. The sign on the road says to turn right. Overall, the final answer is ’turn right’.
Subgoal-based plan:
Subgoal 1: Describe the shape of the sign on the road lane the cameraman is in from the image.
Action 1-1: R1 = VQA([IMG], Question: What’s the sign on the road lane?) = There’s a right turn arrow on the road.
Subgoal 2: Answer which lane the cameraman is in.
Action 2-1: R2 = QA([R1], Question: Which lane is the cameraman in?) = The cameraman is in right turn lane.
Subgoal 3: Answer what the cameraman has to do from his position.
Action 3-1: R3 = QA([R1, R2], Question: What does the cameraman have to do from his position?) = turn right.
Subgoal 4: Select the choice from the answer choices.
Action 4-1: R4 = VQA([IMG], Question: If the cameraman were driving what do they have to do from this
position? We can also see that There’s a right turn arrow on the road, The cameraman is in right turn
lane, turn right. There’re some choices: A. turn left, B. drive straight, C. reverse course, D. turn
right. You must choose one from ’A’, ’B’, ’C’ or ’D’. Output only one letter!) = D
@@
Appendix K In-Context Example for Adapting Modules to Unseen Task — InterCodeSQL
Similar to §J, as InterCodeSQL is another unseen task, we showcase one of the three in-context InterCodeSQL-related examples used for the inference of planning and grounding modules, respectively.
K.1 In-Context InterCodeSQL-Related Example for Planning Module
@@
K.2 In-Context InterCodeSQL-Related Example for Grounding Module